A Bibliography of Publications in Byte Magazine: 1970–1979

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

12 January 2024 Version 2.07

Title word cross-reference

Application [WG79]. Approximations [Ruc78]. Arithmetic [Has77].

Based [LFB79, Sto79]. Boolean [WG79].

calculation [Bow79]. Chess [Wha78]. Computer [Wha78]. control [Gor79]. curve [Bow79]. curve-plotting [Bow79].

Design [Taf79]. dimensional [Got78]. Directed [Kor79]. Don’t [All79a].

Education [LFB79].

Floating [Has77]. Functional [Ruc78].

Hidden [Got78].

Interpreter [Taf79]. Invocation [Kor79].

Lambdino [PR79]. Languages [Kor79]. line [Got78]. LISP [All79a, All79b, LFB79, Pra79, Sto79, Taf79, WG79]. Logic [WG79].

M6800 [Taf79]. Management [PR79]. Math [Bon78]. Mathematician [Pra79]. Mathematics [Sto79]. Minimizing [Bow79].

Overlook [All79a]. Overview [All79b].

Pattern [Kor79]. Pattern-Directed [Kor79]. plotting [Bow79, Got78]. Poem [Ste73]. Point [Has77].

Queueing [Gor79].

Real [Bon78].

science [Gor79]. Storage [PR79]. subroutines [Got78]. Symbolic [Sto79]. System [PR79]. Systems [LFB79, Sto79].

theory [Gor79]. three [Got78]. three-dimensional [Got78]. Trees [Ste73]. Tutorial [Wha78].

View [Pra79].

wait [Gor79]. World [Bon78].

References

Allen:1979:DOL

[All79a] J. R. Allen. Don’t overlook LISP.

Byte Magazine, 4(3):??, March 1979. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Allen:1979:OL

[All79b] J. R. Allen. An overview of LISP.

Byte Magazine, 4(8):???–122, August 1979. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Boney:1978:MRW

[Bon78] J. Boney. Math in the real world.

Byte Magazine, 3(9):114–119, September 1978. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bowker:1979:MCP

[Bow79] T. G. Bowker. Minimizing curveplotting calculation. Byte Magazine, 4(12):134–182, December 1979. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gorney:1979:QTS

[Gor79] L. Gorney. Queueing theory, the science of wait control. Byte Magazine, 4(4, 5):??, ?? 1979. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gottlieb:1978:HLS

[Got78] M. Gottlieb. Hidden line subroutines for three-dimensional plotting. Byte Magazine, 3(5):49–58, May 1978. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hashizume:1977:FPA

[Has77] B. Hashizume. Floating point arithmetic. Byte Magazine, 2(11):76–78, 180–188, November 1977. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kornfeld:1979:PDI

[Kor79] W. A. Kornfeld. Pattern-directed invocation languages. Byte Magazine, 4(8):34–48, August 1979. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Laubsch:1979:LBS

[LFB79] Joachim H. Laubsch, G. Fischer, and H. D. Bocker. LISP based systems for education. Byte Magazine, 4(8): 18–24, August 1979. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Prini:1979:LSM

[PR79] G. Prini and M. Rudalics. The Lambdino Storage Management System.

Byte Magazine, 4(8):26–32, August 1979. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pratt:1979:MVL

[Pra79] V. R. Pratt. A mathematician’s view of LISP. Byte Magazine, 4(8):162– 168, August 1979. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ruckdeschel:1978:FA

[Ruc78] F. Ruckdeschel. Functional approximations. Byte Magazine, 3(11):34– 46, November 1978. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Steele:1973:TP

[Ste73] Guy Lewis Steele Jr. Trees (poem).

Byte Magazine, ??(??):??, ?? 1973. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Stoutemyer:1979:LBS

[Sto79] D. R. Stoutemyer. LISP based symbolic mathematics systems. Byte Magazine, 4(8):176–192, August 1979. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Taft:1979:DML

[Taf79] S. T. Taft. The design of an M6800 LISP interpreter. Byte Magazine, 4 (8):132–152, August 1979. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Weyhrauch:1979:LAB

[WG79] R. Weyhrauch and H. Graves. LISP application in Boolean logic. Byte Magazine, 4(8):206–211, August 1979.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Whaland:1978:CCT

[Wha78] Norman Whaland. A computer chess tutorial. Byte Magazine, 3(10):168– ??, October 1978. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).